OpenStudy (anonymous):

Find the square. (2x – 6)2

OpenStudy (paxpolaris):

\[(2x – 6)^2=\left( 2x-6 \right)\left( 2x-6 \right)\]

OpenStudy (paxpolaris):

\[\large=2x(2x-6)-6(2x-6)\]